Calculate the hydronium ion concentration and the pH when 15.0 mL of 0.80 M  is mixed with 15.0 mL of 0.80 M  ().

Concentration = M

pH =
Answer #1

if the two given solutions are acidic,

number of moles of one solution = 0.80M x 0.015L = 0.012 mol

number of moles of second solution = 0.80M x 0.015L = 0.012 mol

Total volume of the solution = 15.0+15.0= 30.0 mL = 0.030L

after mixing

concentration of resultant solution = 0.012 + 0.012 / 0.030

[H+] = 0.8 M

Concentration of H3O+ ion = 0.8M

[H+] = 0.8M

-log[H+] = -log(0.8)

PH = 0.097